OpenStudy (anonymous):

Use the formula y = 1.5x + 400 to predict the rice weevil population 10 weeks after the beginning of the study.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

A variable is a symbol used to represent one or more numbers. Replace the letter "x" with the number 10 and simplify to find what y is for that value of x.

OpenStudy (skullpatrol):

$$\Huge y = 1.5x + 400 $$ $$\Huge y = 1.5(10) + 400$$

OpenStudy (anonymous):

1.5(10) ? not to the power of ten ?

OpenStudy (skullpatrol):

$$\Huge y = 1.5x + 400 = 1.5*x+400$$

OpenStudy (skullpatrol):

the power of ten would look like this: $$ y = 1.5^x + 400$$

OpenStudy (skullpatrol):

$$\Huge y = 1.5^x + 400 $$